# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ 2.2 Differential Privacy Definitions
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{O(r\log(dR/(\epsilon_{s}p_{f})))}:\textsc{DPTreeSoftmax}$triangleright Algorithm 3
@@
6:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
7:     for $i=1\to l$ do
8:         ${\mathcal{D}}_{i}$.Init$(X,n,w,\epsilon/l,\delta/l,\delta^{\prime}/l,c,\epsilon_{s})$
@@
10:end procedure
11:procedure DistanceQuery($y\in[0,R]^{d},\alpha\in(0,1)$)
12:     $l\leftarrow O(r\log(dR/(\epsilon_{s}p_{f})))$
13:     $r\leftarrow 0^{l}$
14:     for $i=1\to l$ do
15:         $r_{i}\leftarrow{\mathcal{D}}_{i}.\textsc{DistanceQuery}(y,\alpha)$
@@
17:     return Median of $r$
18:end procedure
19:end datastructure
3 Main Results: Cross-Attention
In this section, we show our main result for cross-attention. Theorem 3.1 states that we can ensure the entire cross-attention module satisfies DP and is robust to adaptive queries. Our high-level idea is based on the similarity between weighted distance problem and cross-attention. For a typical weighted distance problem, we define the following: Let $w\in\mathbb{R}^{n}$ be the weights, $X\in\mathbb{R}^{n\times d}$ be the data matrix, where $x_{i}^{\top}$ is the $i$-th row of $X$ for $i\in[n]$, and let $y\in\mathbb{R}^{d}$ be the query. Suppose we need to answer $\ell_{1}$-distance query. We have
@@ 3 Main Results: Cross-Attention
which can be viewed as a weighted Softmax problem, where $V$ provides the weights, $Q$ is the query, and $K$ is the dataset. Thus, we choose to add noise to $K$ and $V$ based on the similarity between the weighted distance problem and cross-attention. Furthermore, we find that we can only handle one column of $V$, i.e., $V_{*,k}\in\mathbb{R}^{n}$, in a single data structure. Therefore, we need to initialize a total of $d$ different data structures, each with weights $V_{*,k}$ for $k\in[d]$.
Here, we present our main result below.
Theorem 3.1 (Softmax cross-attention, informal version of Theorem I.12).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be the parameters of the polynomial kernel methods (Lemma C.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention satisfies DP and supports the following operations:
  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 1), for $k\in[d]$. It takes $O(lnr)$ time to initialize one data structure.
  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 1) for each $i\in[m]$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
In Theorem 3.1, we use our DPTreeSoftmaxAdaptive (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta)$-DP with $n^{-1}(\alpha+\epsilon_{s})$ relative error and $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(\alpha^{-1}lr\log^{2}n)$ running time under adaptive query. More specifically, the algorithm creates $d$ DPTreeSoftmaxAdaptive data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time. Notably, our error is inversely proportional to $n$, meaning that as the input token length increases, both the relative and approximate errors approach zero. This is achieved by the normalizing matrix $D$ (Definition 1.1). We refer the reader to Section I for proof details.
Algorithm 2 DPTree initialization and query
1:datastructure DPTree $\triangleright$ Theorem C.1
2:members
3:     $b:\mathbb{R}^{2n-1}$
4:     $c:\mathbb{R}^{2n-1}$
5:end members
6:procedure Init($a\in\mathbb{R}^{n},n\in\mathbb{N}_{+},\Delta\in\mathbb{R},\epsilon\in(0,1),\delta\in(0,1)$) $\triangleright$ Lemma D.3, Lemma C.3
7:     $b[n,2n-1]\leftarrow a$
8:     for $i=n\to 2n-1$ do
9:         $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
10:     end for
11:     for $i=(\log n)\to 1$ do
12:         for $j=1\to 2^{i-1}$ do
13:              $k\leftarrow 2^{i-1}+j-1$
14:              $b[k]\leftarrow b[2k]+b[2k+1]$
15:              $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
16:         end for
17:     end for
18:end procedure
19:procedure Query($x\in[n],y\in[n]$) $\triangleright$ Lemma D.4, D.5, D.6
20:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $c$) by using at most $2\log n$ nodes on the path. Let Value be the above summation.
21:     return Value
22:end procedure
23:procedure TrueQuery($x\in[n],y\in[n]$)
24:     Trace from bottom nodes of $x$ and $y$ to find their lowest common ancestor, then we report the summation (based on $b$) by using at most $2\log n$ nodes on the path, where the height of the tree is $\log n$, and we need left and right boundary points. Let Value be the above summation.
25:     return Value
26:end procedure
27:end datastructure
@@ 4 Key Data Structure: DPTree
This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 5 and 6), DPTreeHighDim (Algorithm 7), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 1).
In Section 4.1, we provide our high-level proof insights. In Section 4.2, we give our basic building block algorithms DPTree, DPTreeDistance and DPTreeHighDim. In Section 4.3, we present our DPTreeSoftmax algorithm that solves the weighted Softmax problem. In Section 4.4, we present our DPTreeSoftmaxAdaptive algorithm that enables DPTreeSoftmax to handle adaptive query problem.
@@ 4.1 Technique Overview
Notice that Eq. (1) is not a typical distance measure like $\ell_{1}$ or $\ell_{2}$, but by using polynomial kernel method techniques, we transform it into a distance measure. AS [6] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error. Then, by the Law of Cosines, we transform the inner product of polynomial kernel functions into a distance measure, i.e.,
$2P(x)^{\top}P(y)=$ $~{}-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide cross-attention DP guarantee.
In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the leaf nodes to ensure DP guarantee. Then, based on this data structure, we design DPTreeDistance (Algorithm 5 and 6) to answer one dimensional weighted distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|$, which utilizes DPTree to store and return noised weights $w_{i}$ multiplied with the approximated distances between the query $y$ and data $x_{i}$. We further decompose high dimensional $\ell_{p}^{p}$-distance problem into one dimensional $\ell_{1}$-distance problems using
@@
Based on this decomposition, we design DPTreeHighDim (Algorithm 7) which is capable of answering high dimension queries. Then, using Eq. (2) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries. By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeSoftmaxAdaptive (Algorithm 1). See Section C for a more detailed outline of algorithms and proof techniques.
@@ 4.2 DPTree, DPTreeDistance, and DPTreeHighDim
We design a basic data structure DPTree (Algorithm 2) that answers summation queries by a summation segment tree with truncated Laplace noise (Definition 2.4). The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node. In query time, we first trace from bottom nodes to find their lowest common ancestor, then report the summation by using at most $2\log n$ nodes on the path (Algorithm 2). Based on the parallel composition rule of DP (Fact B.7), we find that if we have multiple disjoint interval queries, the error of the weighted sum of the intervals can be bounded independently of the number of queries (Lemma D.8). See more details in Section D.
We then design DPTreeDistance, a one-dimensional weighted $\ell_{1}$ distance data structure detailed in Algorithm 5 and 6. Initialization involves rounding each data point to the nearest multiple of a small interval and aggregating their weights into an array (illustrated in Figure 1), which is then input into our DPTree. At query time, we retrieve aggregated weights within small intervals and multiply them by their distances to the query point. We introduce a relative error parameter $\alpha$ to reduce the number of iterations to $O(\log(n)/\alpha)$, improving efficiency. Guided by Eq.(3), we design DPTreeHighDim (Algorithm 7), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate. See details in Section F and G.
@@
3:     ${\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$triangleright Algorithm 5, Theorem I.7
4:     $P:[0,\Gamma_{R,s}]^{n\times r}$triangleright Definition I.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
@@
8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma C.7
@@
11:         Compute $P(x_{j})$triangleright Polynomial kernel function $P(\cdot)$, Lemma I.5
12:         Compute $w_{j}\|P(x_{j})\|_{2}^{2}$
13:         $P_{wx}\leftarrow P_{wx}+w_{j}\|P(x_{j})\|_{2}^{2}$
14:         $s_{w}\leftarrow s_{w}+w_{j}$
15:         $P_{j,:}\leftarrow P(x_{j})$
16:     end for
17:     for $i=1\to r$ do
18:         $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $c\epsilon/\sqrt{r\log(1/\delta^{\prime})}$, $\delta/r$) $\triangleright$ Algorithm 5
19:     end for
@@
21:procedure DistanceQuery($y\in[0,R]^{d}$, $\alpha\in(0,1)$) $\triangleright$ Lemma C.7
@@
26:         Value $\leftarrow$ Value + $\mathcal{D}_{i}$.DistanceQuery$(P(y)_{i},\alpha)$triangleright Algorithm 6

### v3

@@ 4 Main Results: Cross-Attention
3:   ${\mathcal{D}}_{0},{\mathcal{D}}_{1},\dots,{\mathcal{D}}_{d}:\textsc{DPTreeSoftmaxAdaptive}$ $\triangleright$ Algorithm 7
4:end members
5:procedure Init($K\in[0,R]^{n\times d}$, $V\in[-R_{w},R_{w}]^{n\times d}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01)$) $\triangleright$ $n=|K|$
6:  for $k=1\to d$ do
[…]
8:  end for
[…]
10:end procedure
11:procedure Query($Q_{i}\in[0,R]^{d}$)
12:  $O\leftarrow 0^{d}$
[…]
14:  for $k=1\to d$ do
[…]
16:  end for
17:  return $O$
18:end procedure
19:end datastrucutre
Now, we introduce cross-attention. Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. In a standard cross-attention process, $K$ and $V$ are accessible before inference, while the user input $Q$ becomes available only when the user provides it. Here, $K$ and $V$ represent values stored in memory or disks and are considered private assets protected within the model, whereas $Q$ is treated as public.
For the cross-attention mechanism $\mathrm{Attn}$ (Definition 1.1), we aim to ensure that the matrix $AV$ satisfies DP guarantee. Let $A_{i,j}=\exp(\langle Q_{i},K_{j}\rangle/d)$ for $i\in[m],j\in[n]$. Let $V_{j,k}\in\mathbb{R}$ be the $(j,k)$-th entry of $V$, for $j\in[n],k\in[d]$.
The $(i,k)$-th entry of $AV$ for each $i\in[m],k\in[d]$ is computed by
which can be viewed as a weighted Softmax problem, where $V$ provides the weights, $Q$ is the query, and $K$ is the dataset. Thus, we choose to add noise to $K$ and $V$ based on the similarity between the weighted distance problem and cross-attention. Furthermore, we find that we can only handle one column of $V$, i.e., $V_{*,k}\in\mathbb{R}^{n}$, in a single data structure. Therefore, we need to initialize a total of $d$ different data structures, each with weights $V_{*,k}$ for $k\in[d]$. For computing $D$, we treat $V={\bf 1}_{n}$, which can be interpreted as a weighted Softmax problem with weight ${\bf 1}_{n}$.
Here, we present our main result below.
Theorem 4.1 (Softmax cross-attention, informal version of Theorem G.11).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma G.6). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. Assume the input context length $n$ is large enough. Then there is a data structure DPTreeCrossAttention (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention DP and supports the following operations:
[…]
    • the process of output $z$ has relative error $2\epsilon_{s}/(1-\epsilon_{s})$ and additive error $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n),$
    • it holds with probability $1-p_{f}$ (where $p_{f}$ is used in $l$),
    • it is robust to adaptive query.
Remark 4.2.
Notice in Theorem 4.1 that we ensure the process of computing each entry is $(\epsilon,\delta+\delta^{\prime})$-DP. To guarantee that the overall output vector of length $d$ is DP, we initialize each $\mathcal{D}_{i}$ for $i\in\{0,1,2,\dots,d\}$ with parameters scaled from $\epsilon/2,\delta/2,\delta^{\prime}/2$ to $\epsilon/(d+1),\delta/(d+1),\delta^{\prime}/(d+1)$. Then, by the basic composition property (Fact A.8), the output vector is $(\epsilon,\delta+\delta^{\prime})$-DP, with the additive error increasing by a factor of $\widetilde{O}(d)$.
In Theorem 4.1, we use our DPTreeCrossAttention (Algorithm 1) and guarantee that, for each query token of cross-attention, the output process satisfies $(\epsilon,\delta+\delta^{\prime})$-DP with $2\epsilon_{s}/(1-\epsilon_{s})$ relative error and $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$ additive error, and $O(ldr\log n)$ running time under adaptive query. More specifically, the algorithm creates $d+1$ DPTreeSoftmaxAdaptive (Algorithm 7) data structures, each requiring $O(lnr)$ memory consumption and $O(lnr)$ initialization time. Notably, our additive error is inversely proportional to $n$, meaning that as the input token length increases, the additive error approaches zero. We refer the reader to Section G for proof details.
Algorithm 2 DPTree initialization and query
1:datastructure DPTree $\triangleright$ Theorem B.1
2:members
3:   $c:\mathbb{R}^{2n-1}$
@@
6:  $b[n,2n-1]\leftarrow a$
7:  for $i=n\to 2n-1$ do
8:   $c[i]\leftarrow b[i]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
@@
10:  for $i=(\log n)\to 1$ do
11:   for $j=1\to 2^{i-1}$ do
12:     $k\leftarrow 2^{i-1}+j-1$
13:     $b[k]\leftarrow b[2k]+b[2k+1]$
14:     $c[k]\leftarrow b[k]+\mathrm{TLap}(\Delta,\epsilon/\log n,\delta/\log n)$
15:   end for
@@
17:end procedure
18:procedure Query($y\in[0,R]$)
19:  $c_{\rm left},c_{\rm right}\leftarrow 0,0$
20:  for $i=1\to\log n$ do
21:   Let node $j\in[2^{i}]$ of layer $i$ denotes the integer such that $y\in[(j-1)R/2^{i},jR/2^{i})$
22:   if $j$ is even then $\triangleright$ Node $j$ is the right child of its parent
23:     $c_{\rm left}\leftarrow c_{\rm left}+c[2^{i}+j-2]$ $\triangleright$ Add the value of left sibling node
@@
25:     $c_{\rm right}\leftarrow c_{\rm right}+c[2^{i}+j]$ $\triangleright$ Add the value of right sibling node
26:   end if
27:  end for
28:  return $c_{\rm left},c_{\rm right}$
29:end procedure
30:end datastructure
@@ 5 Key Data Structure: DPTree
This section provides our key data structures: DPTree (Algorithm 2), DPTreeDistance (Algorithm 4 and 5), DPTreeHighDim (Algorithm 6), DPTreeSoftmax (Algorithm 3), and DPTreeSoftmaxAdaptive (Algorithm 7).
@@ 5.1 Technique Overview
Notice that Eq. (1) is not a typical distance measure like $\ell_{1}$ or $\ell_{2}$, but by using polynomial kernel method techniques, we transform it into a distance measure. [4] states that the exponential inner product can be approximated by polynomial kernel function $P(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{r}$, i.e., $P(x)^{\top}P(y)\approx\exp(x^{\top}y/d)$ for two vector $x,y\in\mathbb{R}^{d}$, with a relative error. Then, by the Law of Cosines, we transform the inner product of polynomial kernel functions into a distance measure, i.e.,
$~2P(x)^{\top}P(y)$
$=$ $~-\|P(x)-P(y)\|_{2}^{2}+\|P(x)\|^{2}_{2}+\|P(y)\|^{2}_{2}.$ (2)
After transforming Eq. (1) into a distance measure, we design the DPTree series data structures to provide a cross-attention DP guarantee.
In summary, we first design the data structure DPTree (Algorithm 2) that builds a binary segment tree with truncated Laplace noise added in the nodes to ensure DP guarantee. Then, based on this data structure, we design DPTreeDistance (Algorithm 4 and 5) to answer one dimensional weighted $\ell_{p}^{p}$ distance queries $\sum_{i=1}^{n}w_{i}\cdot|y-x_{i}|^{p}$. We further decompose high dimensional $\ell_{p}^{p}$ distance problem into one dimensional $\ell_{p}^{p}$ distance problems using
@@
Based on this decomposition, we design DPTreeHighDim (Algorithm 6), which is capable of answering high-dimension queries. Then, using Eq. (5.1) and DPTreeHighDim, we design DPTreeSoftmax (Algorithm 3) to answer Softmax queries. By building multiple copies of this data structure, we boost the success probability such that it can answer any query (including adaptive query) with an additive error, establishing the final data structure DPTreeCrossAttention (Algorithm 1).
@@ 5.2 DPTree, DPTreeDistance, and DPTreeHighDim
The unweighted distance query has been explored in prior works [36, 8, 45]. Specifically, [36] leverages online learning techniques to approximate the sum of distances, while [8] introduces a DP data structure based on a node-contaminated balanced binary tree. Furthermore, [45] presents a new data representation in tree nodes, where each node stores the sum of distances from one point to multiple points. In contrast, we focus on the weighted distance query, generalizing their results.
We design a basic data structure DPTree (Algorithm 2) that answers summation queries by a summation segment tree with truncated Laplace noise (Definition 3.4). The algorithm first builds a binary summation tree in an array and then adds truncated Laplace noises to each node.
During a query, the algorithm traverses each layer of the binary structure based on the input $y$, aggregating values from sibling nodes by accessing at most $O(\log n)$ nodes along the path. It then returns the accumulated left and right sums as the query result (Algorithm 2). See more details in Section B.
We then design DPTreeDistance, a one-dimensional weighted $\ell_{p}^{p}$ distance data structure detailed in Algorithm 4 and 5.
Initialization involves assigning each data point to the nearest bin and aggregating their weighted polynomial terms into multiple arrays (illustrated in Figure 1), which are then used to initialize several instances of our DPTree. At query time, the algorithm retrieves aggregated weights from each DPTree corresponding to the query point and combines them using binomial coefficients and distance powers to compute the one-dimensional weighted $\ell_{p}^{p}$ distance. Guided by Eq. (3), we design DPTreeHighDim (Algorithm 6), which extends DPTreeDistance to higher dimension by constructing independent data structures for each coordinate. See details in Section D and E.
@@
3:   ${\mathcal{D}}_{0},{\mathcal{D}}_{1},\dots,{\mathcal{D}}_{r}:\textsc{DPTreeDistance}$ $\triangleright$ Algorithm 4, Theorem D.1
4:   $P:[0,\Gamma_{R,s}]^{n\times r}$ $\triangleright$ Definition G.3 for $\Gamma_{R,s}$, Eq. (9) for $s$, Eq. (10) for $r$
@@
8:procedure Init($X\subset[0,R]^{d}$, $n\in\mathbb{N}_{+}$, $w\in[-R_{w},R_{w}]^{n}$, $\epsilon\in(0,1)$, $\delta\in(0,1)$, $\delta^{\prime}\in(0,1)$, $c\in(0,0.1)$, $\epsilon_{s}\in(0,0.1)$) $\triangleright$ Lemma G.6
@@
11:   Compute $P(x_{j})$ $\triangleright$ Polynomial kernel function $P(\cdot)$, Lemma G.5
12:   $P_{j,:}\leftarrow P(x_{j})$
13:  end for
14:  for $i=1\to r$ do
15:   $\mathcal{D}_{i}.$Init($P_{:,i}$, $n$, $w$, $\frac{c\epsilon}{3\sqrt{r\log(2/\delta^{\prime})}}$, $\frac{\delta}{3r}$) $\triangleright$ Algorithm 4
16:   $P_{wx}\leftarrow P_{wx}+\mathcal{D}_{i}$.DistanceQuery(0)(0)
17:  end for
18:  $\mathcal{D}_{0}.$Init(${\bf 1}_{n}$, $n$, $w$, $\epsilon/3$, $\delta/3$)
19:  $s_{w}\leftarrow s_{w}+\mathcal{D}_{0}$.DistanceQuery(0)(0)
@@
21:procedure DistanceQuery($y\in[0,R]^{d}$) $\triangleright$ Lemma G.6
@@
[…]